OpenStudy (anonymous):

In the list of 3 consecutive integers, the sum of the smallest and twice the largest is 134. what is the smallest of the 3 integers?

OpenStudy (johnweldon1993):

Let the first number = x The second = x + 1 The third = x + 2 So we know the sum of the smallest and twice the largest = 134 \[\large x + 2(x + 2) = 134\] Now we can solve that , first simplify by distributing that 2 through the parenthesis \[\large x + 2x + 4 = 134\] Can you take it from there?

OpenStudy (anonymous):

x = 130/3 and is not an integer.

OpenStudy (anonymous):

@sahi3 Are you sure the problem statement is correct ?

OpenStudy (anonymous):

i forgot to put 3 comsecutive even numbers

OpenStudy (anonymous):

Solve the following or x: x + 2 (x + 4) = 134

OpenStudy (anonymous):

why is it x+4

OpenStudy (anonymous):

Let x be the smallest even number. Then the next two consecutive even numbers larger than x is x+2 and x+4 respectively.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

You are welcome.